OpenStudy (anonymous):

The midpoint of a line segment with end points as (2, 3) and (4, y2) is (3, 2). What is the value of y2? -2 -1 1 7

OpenStudy (cwrw238):

midpoint = (2+4)/2, (3+y2)/ 2 = (3,(3+y2)/ 2 (3+y2)/ 2 = 2 3+y2 = 4 y2 = 1

OpenStudy (anonymous):

Thanks SO much! I really appreciate the help. Geometry and I don't quite get along.

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

y2 = 1

OpenStudy (anonymous):

Thanks!